OpenStudy (anonymous):

U-Substitution: Evaluate integral of (20x^4/(4x^5 + 3))dx; u=4x^5 + 3 I found dx but I'm not sure if the answer is integral of (1/(4x^5 + 3))

OpenStudy (astrophysics):

\[u = 4x^5+3\] then we have \[\frac{ du }{ dx } = 20x^4 \implies du = 20x^4dx\] now compare it with your integrand, plug it all in.

OpenStudy (anonymous):

yes, then dx = du/20x^4 and you would have integral of (20x^4/u)(du/20x^4) do i cancel out both 20x^4?

OpenStudy (astrophysics):

No, we don't solve for dx when we make the substitution, we are solving for du, since we want to simplify the integrand. Compare the integral with the substitution \[\int\limits \frac{ \color{red}{20x^4dx} }{ 4x^5+3 }\]

OpenStudy (astrophysics):

We replace everything in the integral in terms of u

OpenStudy (anonymous):

okay, always solve for du and not dx?

OpenStudy (astrophysics):

Yes, if you're given a integral as such.

OpenStudy (anonymous):

The denominator becomes u

OpenStudy (astrophysics):

Right

OpenStudy (astrophysics):

How about the numerator what happens

OpenStudy (anonymous):

you find the integral of the numerator? 4x^5?

OpenStudy (astrophysics):

Ok lets take this slow, so we have our integral \[\int\limits \frac{ 20x^4 }{ 4x^5+3 }dx\] we start by noticing u - substitution is the best method as the degree in the denominator is a degree higher so it will work nicely. Then we make our u substitution \[u=4x^5+3\]\[\frac{ du }{ dx } = 20x^4 \implies du = \color{red}{20x^4dx}\] now we look back at our integral and compare, \[\int\limits \frac{ \color{red}{20x^4} }{ 4x^5+3 }\color{red}{dx}\] notice our du is exactly the same as when we did the derivative for du/dx. So plugging it all in since we are making it in terms of u we get, \[\int\limits \frac{ \color{blue}{du}}{ u }\] notice our integral has been simplified now since all we have to do is make our "substitution in terms of u."

OpenStudy (anonymous):

okay, so its du/4x^5 + 3?

OpenStudy (astrophysics):

What do you mean

OpenStudy (astrophysics):

\[u=4x^5+3\] So we have \[\int\limits \frac{ du }{ u }\]

OpenStudy (anonymous):

u = 4x^5 + 3 and if you plug that into du/u is it du/4x^5 + 3?

OpenStudy (astrophysics):

Nooo, read what I've said, we make the substitution in terms of u, so it's easier to integrate, we will plug our 4x^5+3 after the integration.

OpenStudy (astrophysics):

Now we just go ahead and integrate \[\int\limits \frac{ du }{ u }\]

OpenStudy (anonymous):

okay, how do you integrate that?

OpenStudy (astrophysics):

Note that \[\int\limits \frac{ 1 }{ x } dx = \ln(x)+C\]

OpenStudy (anonymous):

I figured ln had something to do with it. What is next?

OpenStudy (astrophysics):

Please try it yourself, see what you get

OpenStudy (anonymous):

ln(4x^5 + 3)

OpenStudy (astrophysics):

Exactly, don't forget the +C :)

OpenStudy (anonymous):

reeeeaaaaaallly?! wow, thank you! Can you help me with another please?

OpenStudy (astrophysics):

Sure, I'll try

OpenStudy (anonymous):

Same directions \[\int\limits 36x^2e^(4x^3 + 3) dx\] The (4x^3 + 3) is raised

OpenStudy (astrophysics):

Cool, so any idea where to begin? I will let you do most of it this time as it's good practice.

OpenStudy (anonymous):

u = 4x^3 + 3 du = 12x^2dx

OpenStudy (astrophysics):

Good, what's next

OpenStudy (anonymous):

(36x^2e^u)dx

OpenStudy (astrophysics):

Mhm, \[\int\limits 36x^2e^{4x^3+3}dx\] \[u = 4x^3+3 \implies du = 12x^2dx \implies \frac{ du }{ 12 }=x^2dx\] again compare this with your original integral and make all the substitutions correctly.

OpenStudy (anonymous):

why do you divide the 12?

OpenStudy (astrophysics):

Is there a 12 in the integrand originally?

OpenStudy (astrophysics):

We are replacing what already exists

OpenStudy (anonymous):

so because there was no 12 in the original but a x^2, you have to divide by the 12 so the x^2 is alone?

OpenStudy (astrophysics):

Right, because we are "substituting" in for what already exists, if you remember high school math when we had substitution for systems of equations you were replacing equations with substitution.

OpenStudy (astrophysics):

A lot of the math goes back to what you learnt before, I always say the calculus part is easy it's the math you've learnt before that you need to remember.

OpenStudy (anonymous):

okay. so is 36x^2e^(u) correct or is 36(u)e^(4x^5 + 3) correct since we already have x^2?

OpenStudy (anonymous):

I know, I understand most parts of calculus but i always have trouble remembering the algebra

OpenStudy (astrophysics):

It seems you're still having a hard time grasping what substitution means, we don't want the integrand to be a product, that would make it problematic for you. We don't want something like x*e^x this is bad for now, but your integral is set up that requires a "substitution" which gets rid of the product and makes it easy to integrate.

OpenStudy (anonymous):

so what do i have to do after I find that x^2dx = du/12?

OpenStudy (astrophysics):